# Stealing That Free Lunch: Exposing the Limits of Dyna-Style Reinforcement Learning

### v2

@@ Abstract
Dyna-style off-policy model-based reinforcement learning (DMBRL) algorithms are a family of techniques for generating synthetic state transition data and thereby enhancing the sample efficiency of off-policy RL algorithms. This paper identifies and investigates a surprising performance gap observed when applying DMBRL algorithms across different benchmark environments with proprioceptive observations. We show that, while DMBRL algorithms perform well in OpenAI Gym, their performance can drop significantly in DeepMind Control Suite (DMC), even though these settings offer similar tasks and identical physics backends. Modern techniques designed to address several key issues that arise in these settings do not provide a consistent improvement across all environments, and overall our results show that adding synthetic rollouts to the training process — the backbone of Dyna-style algorithms — significantly degrades performance across most DMC environments. Our findings contribute to a deeper understanding of several fundamental challenges in model-based RL and show that, like many optimization fields, there is no free lunch when evaluating performance across diverse benchmarks in RL.
@@ 1 Introduction
Colloquially, the “no free lunch theorem” states that no optimization algorithm can be universally optimal across all problem instances. In reinforcement learning (RL), this implies that performance will vary depending on the environment and problem characteristics. In that light, this paper begins with a simple, yet novel and unexpected observation: Model-Based Policy Optimization (MBPO) (Janner et al., 2019), a popular Dyna-style (Sutton, 1991) model-based reinforcement learning (DMBRL) algorithm, demonstrates strong performance across tasks in OpenAI Gym (Brockman et al., 2016), but performs significantly worse than its base off-policy algorithm, Soft Actor Critic (SAC) (Haarnoja et al., 2019b), when trained in DeepMind Control Suite (DMC) (Tassa et al., 2020) — cf. Figure 1.
Interestingly, the Gym and DMC benchmarks feature similar tasks and identical physics backends (Todorov et al., 2012). Yet, despite the popularity of MBPO, with around $1$k citations (as of late 2024), only two studies report evaluating MBPO in DMC (Wang et al., 2023a; Voelcker et al., 2024), and a performance gap has only been noted in the hopper-hop task (Voelcker et al., 2024). Our results generalize these findings, and call into question the robustness of MBPO, and of Dyna-style algorithms more broadly.
@@
Figure 1: Comparison of normalized final return for two different DMBRL algorithms with and without Dyna-style enhancements. Results are averaged across 6 random seeds per task, with 6 tasks from OpenAI Gym (Brockman et al., 2016) and 15 from DMC (Tassa et al., 2020). For training curves, cf. Figures 3 and 15.
@@
Our extensive results indicate: “yes they are fundamental, and no we cannot.” Figure 1 shows that this gap persists more broadly than in just MBPO, but also in another recent DMBRL algorithm, Aligned Latent Models (ALM). These algorithms have contrasting design philosophies: MBPO is built on SAC (Haarnoja et al., 2019b), a stochastic policy algorithm leveraging entropy-based exploration, while ALM is built on DDPG (Lillicrap et al., 2019), a deterministic policy algorithm that relies on additive Gaussian noise for exploration. These differences in exploration strategies, combined with their distinct architectures and objectives, make MBPO and ALM complementary test cases for assessing whether the observed gap is inherent to Dyna-based methods rather than specific to a particular implementation.
@@
Using MBPO as a surrograte for a subclass of DMBRL algorithms that underperform in DMC, we analyze the Gym/DMC performance gap and try to close it by addressing some possible sources of this discrepancy including overestimation bias, neural network plasticity, and environment modeling fidelity. Despite these efforts, there remains a substantial gap in performance on these two benchmarks, which suggests that the generalization challenges of MBPO — and, by extension, a subclass of DMBRL algorithms — are more deeply rooted.
@@
Second, the fact that this oversight has remained unnoticed for several years highlights a significant, pervasive issue which has also noted in several other AI subfields (Haibe-Kains et al., 2020; Bender et al., 2021; Agarwal et al., 2022; Ahmed et al., 2023; Jordan et al., 2024): a lack of reasonable access which prevents scientists from reproducing and verifying key results. For instance, per the calculations in Appendix B, generating just Figure 4 of this manuscript using the original MBPO implementation (Janner, 2019) would require over $100$ GPU-days, assuming optimistic runtime estimates from prior work (Ghugare et al., 2023; Xu et al., 2022).
@@
Such computational barriers limit experimentation and cross-comparison, leaving critical issues, like those we highlight, unexamined. To address this, we developed a new MBPO implementation built on a high-performance SAC implementation (D’Oro et al., 2023), leveraging JAX (Bradbury et al., 2018) for efficient parallelization. This new implementation reduces training time for the experiments in Figure 4 to approximately $4$ GPU-days, a 26×26$\times$26 × speedup compared to the original Pytorch implementation (Janner et al., 2019).
@@
Our implementation also compares favorably to Aligned Latent Models (ALM), a model-based algorithm known for its high wall-clock speed, as shown in Figure 2. These combined improvements over previous model-based RL implementations empower researchers with limited computational resources to conduct extensive DMBRL studies using just a single GPU.
@@
Figure 2: Comparison of seconds per environment step across multiple DMBRL implementations when deployed on 6 OpenAI Gym environments. Compared to Pytorch MBPO and ALM, our implementation takes $\sim$40×40$\times$40 × and $\sim$4×4$\times$4 × less time per environment step, respectively.
@@
    Empirically demonstrating that DMBRL methods can suffer from a significant performance gap when training from scratch in OpenAI Gym versus DMC environments.
@@
    Dramatically accelerating the DMBRL experimentation process with a new JAX-based implementation, which achieves up to a $\sim$40×40$\times$40 × decrease in wall-clock time. This acceleration lowers the computational barrier for researchers to develop (and comprehensively evaluate) DMBRL algorithms.
@@ 2.1 Reinforcement Learning, Model-Based RL, and Dyna-style Algorithms
Like most off-policy methods, MBPO maintains a replay buffer of the transitions it has experienced in the environment. MBPO supplements this real world data by generating synthetic on-policy rollouts that branch from states in the replay buffer, thereby augmenting the replay data with “imagined” transitions. This combined, partially synthetic, dataset is then used to train both the actor, a neural network that determines the optimal action based on the current state, and the critic, a neural network that estimates the expected return given a choice of action and state. For MBPO this process is done using the off-the-shelf model-free algorithm Soft Actor-Critic (SAC) (Haarnoja et al., 2019b), a widely used, reliable method for continuous control tasks. MBPO has demonstrated strong performance on OpenAI Gym benchmarks, often greatly surpassing traditional model-free approaches in sample efficiency. Due to its reputation for strong performance, many recent algorithms are based on MBPO, e.g. (Lai et al., 2020; Li et al., 2024; Dong et al., 2024; Zheng et al., 2023; Wang et al., 2023b; Lai et al., 2022).
Another recent extension of the Dyna line of model-based RL is Aligned Latent Models (ALM) (Ghugare et al., 2023). Rather than generating synthetic data in the true state-action space like MBPO, ALM jointly learns observation representations, a world model that predicts next representations given the current representation, and a policy that acts in the representation space. It then uses the Deep Deterministic Policy Gradient (DDPG) algorithm (Lillicrap et al., 2019) for off-policy learning and uses synthetic trajectories to train the latent policy only. Like MBPO, ALM has demonstrated strong performance across OpenAI Gym tasks, while also requiring substantially less wall clock time than MBPO for training, as shown in Figure 2.
Finally, a particularly notable DMBRL approach is DreamerV3 (Hafner et al., 2024), which, unlike the other DMBRL methods presented in this work, was not tested in Gym, but has state-of-the-art sample efficiency in DMC. Like ALM, DreamerV3 optimizes a policy directly in the latent space using imagination-based trajectories and actor-critic learning. Unlike ALM, which uses a single objective for optimizing both the policy and latent model, DreamerV3 separates these processes and learns a latent world model using multiple objectives. Please refer to Appendix A for further details.
@@ 3 Performance Gap of Dyna-based MBRL Across Benchmarks
In this section, we put forth the following hypothesis: incorporating Dyna-style modifications into otherwise successful off-policy algorithms can prevent them from improving beyond their performance at initialization — that of a randomly initialized policy. In support of this hypothesis we present empirical evaluations of MBPO’s performance across six OpenAI Gym tasks and six challenging robotics tasks (Nikishin et al., 2022) from DMC using our JAX-based (Bradbury et al., 2018) implementation of MBPO. For the OpenAI Gym benchmark, we demonstrate that our implementation matches the performance produced by the original PyTorch (Paszke et al., 2019) MBPO implementation (Janner et al., 2019) and ALM (Ghugare et al., 2023). Furthermore, we show that MBPO dramatically outperforms its “no Dyna” base off-policy algorithm in OpenAI Gym, Soft Actor-Critic (SAC) (Haarnoja et al., 2019a), even when both share identical hyperparameters and architectures.
In contrast, in DMC (Tassa et al., 2020) we observe that MBPO struggles to achieve any policy improvements in six out of fifteen [^1] Training curves for MBPO in all fifteen challenging DMC environments are provided in Appendix F in Figure 12. of the challenging environments that we tested on when training from scratch. Because MBPO differs from SAC only due to Dyna-style “enhancements,” these results demand a deeper investigation into whether the default SAC hyperparameters, the model ensemble, or other factors induce an otherwise healthy SAC implementation to fail.
This performance gap extends beyond MBPO into other model-based RL algorithms in the Dyna family (Sutton, 1991). To this end, we deploy the recently developed ALM (Ghugare et al., 2023) method across the same six [^2] Training curves for ALM in all fifteen challenging DMC environments are provided in Appendix F in Figure 15. challenging DMC tasks and observe the same phenomenon.
@@ 3.1 Results in OpenAI Gym
Here, 1-step and $n$-step refer to the length of synthetic rollouts simulated using the learned world model. It is common to prefer shorter rollouts, however, in order to mitigate compounding model errors (Janner et al., 2019; Sikchi et al., 2021); indeed, we demonstrate that these errors are quite substantial in Section 4.1. Combining the results in Figure 2 and Figure 3 we can conclude that our JAX-based implementation of 1-step MBPO matches the original implementation’s (Janner, 2019) $n$-step MBPO’s performance with much greater wall-clock speed, and as such we will henceforth refer to 1-step MBPO as simply “MBPO.”
Figure 3: Performance comparison of our 1-step MBPO implementation, the original MBPO implementation (Janner, 2019), and off-the-shelf implementations of ALM and SAC on OpenAI Gym benchmarks. Solid curves correspond to the mean and shaded regions represent the standard deviation across six trials. The dotted line represents the episodic return of the original $n$-step MBPO at the end of training. 1-step MBPO consistently achieves higher final episodic returns and demonstrates faster convergence compared to SAC. Likewise, 1-step MBPO either matches or surpasses ALM in terms of sample efficiency across all but one task.
@@ 3.2 Results in the DeepMind Control Suite
In contrast, MBPO’s performance when training from scratch drops significantly in the DMC environments despite using the exact same hyperparameters and model structure as for the successes in OpenAI Gym. Full results across all fifteen challenging DMC environments are provided in Figure 12 in Appendix F. These results show the complete range of MBPO’s performance, including scenarios where MBPO solves tasks more slowly than SAC, matches SAC’s sample efficiency, or fails to make any policy improvements.
@@ 3.3 Beyond MBPO: Another Dyna-based Algorithm
In this section, we show that the conclusions in Section 3.2 extend beyond MBPO to another member of the Dyna family, Aligned Latent Models (ALM) (Ghugare et al., 2023). We choose ALM for two reasons. First, unlike MBPO, ALM is built on top of DDPG, which allows us to test Dyna in a deterministic policy framework, in contrast to SAC’s stochastic approach. Second, DDPG’s reliance on adding Gaussian noise to actions during training offers a distinct mechanism for exploration compared to SAC’s entropy-based training strategy.
To evaluate whether Dyna-based “enhancements” negatively impact ALM, we repeated the same experiment as we performed for MBPO across six [^3] Training curves for ALM in all fifteen challenging DMC environments are provided in Appendix F in Figure 15. challenging DMC tasks. In this case we train ALM with and without Dyna-style enhancements using identical hyperparameters, and compare their performance. We note that this “no-Dyna” variant of ALM is inspired by (Ghugare et al., 2023) wherein the authors replaced actor training via synthetic model rollouts with a TD3 (Fujimoto et al., 2018) actor loss that relies only on transitions sampled from the real environment.
@@
Figure 5: Performance of ALM with and without Dyna-style enhancements on six DMC benchmarks. Like MBPO, ALM consistently exhibits no policy improvement across these environments when using Dyna-style enhancements, but exhibits strong performance when they are removed. (6 trials, $\pm$ std).
@@ 4.1 Is High Model Error the Problem?
In Section 3, we saw that Dyna-style “enhancements” caused both MBPO and ALM to dramatically underperform their “no Dyna” counterparts in Deepmind Control (DMC), whereas the same enhancements led to performance gains in OpenAI Gym. Since the predictive model ensemble is central to MBPO’s sample efficiency gains over model-free methods, we focus on this first as a source of MBPO’s issues. Studies have shown that high model error causes significant performance degradation (Gu et al., 2016; Rajeswaran et al., 2017), and models are most effective over short to moderate planning horizons (Sikchi et al., 2021; Holland et al., 2019) if compounding model errors can be mitigated (Talvitie, 2014; Buckman et al., 2019). In this section, we show that the model error for MBPO in many DMC environments is severe enough that even 1-step rollouts are unrealistic, making the pursuit of multi-step methods unproductive.
The relationship between model error and performance degradation in MBPO is challenging to assess, particularly within DMC environments. To investigate this, we compare model errors across six tasks in both OpenAI Gym and DMC. In these experiments, we evaluate the model’s percent error across the training distribution (i.e., the replay buffer) instead of relying on computationally expensive on-policy rollouts. Percent error effectively normalizes across differently-scaled observation spaces, and offers a more intuitive interpretation than other metrics, such as mean squared error.
@@
In Figure 6, we observe that across all six OpenAI Gym tasks, the model error converges significantly below $25\%$. In contrast, in Figure 7, which shows model percent error for the six DMC tasks where MBPO fails, the tasks fall into two categories. The hopper tasks exhibit model errors that converge above $100\%$, while the remaining tasks have model errors converging above $25\%$ as training progresses. Notably, all of these environments fail to improve policy performance beyond that of a random policy (recall Figure 4).
@@
Not all modeling errors present in MBPO when deployed in DMC are significantly larger than those from OpenAI Gym, which makes it difficult to fully contextualize their severity without further insight. To address this gap in understanding, and inspired by previous work (Kalweit & Boedecker, 2017), we next consider tuning the synthetic-to-real data ratio, $S$, for MBPO with a learned model. $S$ represents the proportion of each training batch comprised of synthetic rollouts generated via the learned world model. By varying this ratio, we systematically study the effect of increasing the amount of synthetic data on policy performance and identify how significant the modeling errors are in degrading MBPO’s performance (Gu et al., 2016). Since the default synthetic-to-real data ratio in MBPO is $0.95$, as $S\to 0.95$ we recover MBPO and rely heavily upon synthetic rollouts, whereas as $S\to 0$ we recover SAC, and use only real environment data. We conducted a series of experiments with varying $S$ to evaluate how the fraction of synthetic data impacts the training of the base off-policy algorithm (i.e., SAC).
We focused on two DMC tasks, hopper-stand and humanoid-stand, where MBPO consistently fails to improve policy performance with a learned model. These tasks were chosen because they highlight the limitations of synthetic rollouts in aiding policy improvement and because they cover the two categories of model error magnitude. Our results, summarized in Figure 8, show a clear trend: as the synthetic ratio $S$ increases, episodic return decreases. This effect is particularly evident in Figure 8, where even a small introduction of synthetic data leads to MBPO making no discernible progress on the humanoid-stand task. These findings emphasize that the poor quality of synthetic data for DMC introduces errors that undermine the policy learning component of MBPO.
@@
Since introducing any synthetic data can induce failures in MBPO with the default hyperparameters, we conducted comprehensive hyperparameter sweeps for both hopper-stand and humanoid-stand to rule out tuning-related issues. We swept over the key hyperparameters related to the training and utilization of the predictive model, including the size of the model’s hidden layers, the interval at which the model is retrained, the learning rate, and the number of training steps. None provided any improvement to either model error or policy performance across training, which suggests that deeper architectural and algorithmic changes are required to further reduce model-error. Reduction of model bias remains an open issue in the broader MBRL literature (Wang et al., 2019; Deisenroth & Rasmussen, 2011; Luo et al., 2022) and we hope to address it future work as part of a larger solution to alleviate Dyna-style algorithm issues identified in this work.

### v3

@@ Abstract
Dyna-style off-policy model-based reinforcement learning (DMBRL) algorithms are a family of techniques for generating synthetic state transition data and thereby enhancing the sample efficiency of off-policy RL algorithms. This paper identifies and investigates a surprising performance gap observed when applying DMBRL algorithms across different benchmark environments with proprioceptive observations. We show that, while DMBRL algorithms perform well in control tasks in OpenAI Gym, their performance can drop significantly in DeepMind Control Suite (DMC), even though these settings offer similar tasks and identical physics backends. Modern techniques designed to address several key issues that arise in these settings do not provide a consistent improvement across all environments, and overall our results show that adding synthetic rollouts to the training process — the backbone of Dyna-style algorithms — significantly degrades performance across most DMC environments. Our findings contribute to a deeper understanding of several fundamental challenges in model-based RL and show that, like many optimization fields, there is no free lunch when evaluating performance across diverse benchmarks in RL.
@@ 1 Introduction
Colloquially, the “no free lunch theorem” states that no optimization algorithm can be universally optimal across all problem instances. In reinforcement learning (RL), this implies that performance will vary depending on the environment and problem characteristics. In that light, this paper begins with a simple, yet novel and unexpected observation: Model-Based Policy Optimization (MBPO) (Janner et al., 2019), a popular Dyna-style (Sutton, 1991) model-based reinforcement learning (DMBRL) algorithm, demonstrates strong performance across tasks in OpenAI Gym (Brockman et al., 2016), but performs significantly worse than its base off-policy algorithm, Soft Actor Critic (SAC) (Haarnoja et al., 2019a), when trained in DeepMind Control Suite (DMC) (Tassa et al., 2020) — cf. Figure 1.
Interestingly, the Gym and DMC benchmarks feature similar tasks and identical physics backends (Todorov et al., 2012). Yet, despite the popularity of MBPO, with around $1$k citations (as of late 2024), only two studies report evaluating MBPO in DMC (Wang et al., 2024; Voelcker et al., 2024), and a performance gap has only been noted in the hopper-hop task (Voelcker et al., 2024). Our results generalize these findings, and call into question the robustness of MBPO, and of Dyna-style algorithms more broadly.
Figure 1: Comparison of normalized final return for two different DMBRL algorithms with and without Dyna-style enhancements. Results are averaged across 6 random seeds per task, with 6 tasks from OpenAI Gym (Brockman et al., 2016) and 15 from DMC (Tassa et al., 2020). For training curves, cf. Figures 3 and 17.
@@
Our extensive results indicate: “yes they are fundamental, and no we cannot.” Figure 1 shows that this gap persists more broadly than in just MBPO, but also in another recent DMBRL algorithm, Aligned Latent Models (ALM) (Ghugare et al., 2022). These algorithms have contrasting design philosophies: MBPO is built on SAC (Haarnoja et al., 2019a), a stochastic policy algorithm leveraging entropy-based exploration, while ALM is built on DDPG (Lillicrap et al., 2019), a deterministic policy algorithm that relies on additive Gaussian noise for exploration. These differences in exploration strategies, combined with their distinct architectures and objectives, make MBPO and ALM complementary test cases for assessing whether the observed gap is inherent to Dyna-based methods rather than specific to a particular implementation. While our findings raise serious concerns about these instantiations, our intent is not to challenge the theoretical soundness of the Dyna framework itself, but rather to evaluate how well its practical realizations hold up under realistic, high-fidelity conditions.
@@
Using MBPO as a surrogate for a subclass of DMBRL algorithms that underperform in DMC, we analyze the Gym/DMC performance gap and try to close it by addressing some possible sources of this discrepancy including overestimation bias, neural network plasticity, and environment modeling fidelity. Despite these efforts, there remains a substantial gap in performance on these two benchmarks, which suggests that the generalization challenges of MBPO — and, by extension, a subclass of DMBRL algorithms — are more deeply rooted.
@@
Second, this oversight remaining unnoticed for several years highlights a significant, pervasive issue which has also noted in several other AI subfields (Haibe-Kains et al., 2020; Bender et al., 2021; Agarwal et al., 2021; Ahmed et al., 2023; Jordan et al., 2024): a lack of reasonable access which prevents scientists from reproducing and verifying key results. For instance, per the calculations in Appendix B, generating just Figure 4 of this manuscript using the original MBPO implementation (Janner, 2019) would require over $100$ GPU-days, assuming optimistic runtime estimates from prior work (Ghugare et al., 2022; Xu et al., 2022).
@@
Such computational barriers limit experimentation and cross-comparison, leaving critical issues, like those we highlight, unexamined. To address this, we developed a MBPO implementation built on a high-performance SAC implementation (D’Oro et al., 2023), leveraging JAX (Bradbury et al., 2018) for efficient parallelization. This new implementation reduces training time for the experiments in Figure 4 to approximately $4$ GPU-days, a 26×26$\times$26 × speedup compared to the original Pytorch implementation (Janner et al., 2019).
@@
Our implementation also compares favorably to Aligned Latent Models (ALM), a model-based algorithm known for its high wall-clock speed, as shown in Figure 2. These improvements over previous model-based RL implementations empower researchers with limited resources to conduct extensive DMBRL studies using just a single GPU.
@@
Figure 2: Comparison of seconds per environment step across multiple DMBRL implementations when deployed on 6 OpenAI Gym environments. Compared to Pytorch MBPO and ALM, our implementation takes $\sim$40×40$\times$40 × and $\sim$4×4$\times$4 × less time, respectively.
@@
    Demonstrating that DMBRL methods can suffer from a significant performance gap when training from scratch in OpenAI Gym versus DMC environments.
@@
    Dramatically accelerating the DMBRL experimentation process with a new JAX-based implementation, which achieves up to a $\sim$40×40$\times$40 × decrease in wall-clock time. This acceleration lowers the computational barrier for researchers to develop (and comprehensively evaluate) DMBRL algorithms. We have released our code here: https://github.com/CLeARoboticsLab/STFL.
@@ 2.1 Reinforcement Learning, Model-Based RL, and Dyna-style Algorithms
Like most off-policy methods, MBPO maintains a replay buffer of the transitions it has experienced in the environment. MBPO supplements this real world data by generating synthetic on-policy rollouts that branch from states in the replay buffer, thereby augmenting the replay data with “imagined” transitions. This partially synthetic dataset is then used to train both the actor, a neural network that determines the optimal action based on the current state, and the critic, a neural network that estimates the expected return given a choice of action and state. For MBPO this process is done using the model-free algorithm Soft Actor-Critic (SAC) (Haarnoja et al., 2019a), a widely used, reliable method for continuous control tasks. MBPO has demonstrated strong performance on OpenAI Gym benchmarks, often greatly surpassing traditional model-free approaches in sample efficiency. Due to its reputation for strong performance, many recent algorithms are based on MBPO, e.g. (Lai et al., 2020; Li et al., 2024; Dong et al., 2024; Zheng et al., 2023; Wang et al., 2023; Lai et al., 2021).
Another recent extension of the Dyna line of model-based RL is Aligned Latent Models (ALM) (Ghugare et al., 2022). Rather than generating synthetic data in the true state-action space like MBPO, ALM jointly learns observation representations, a world model that predicts next representations given the current representation, and a policy that acts in the representation space. It then uses the Deep Deterministic Policy Gradient (DDPG) algorithm (Lillicrap et al., 2019) for off-policy learning and uses synthetic trajectories to train the latent policy only. Like MBPO, ALM has demonstrated strong performance across OpenAI Gym tasks, while also requiring substantially less wall clock time than MBPO for training, as shown in Figure 2.
Finally, a particularly notable DMBRL approach is DreamerV3 (Hafner et al., 2025), which, unlike the other DMBRL methods presented in this work, was not tested in Gym, but has state-of-the-art sample efficiency in DMC. Like ALM, DreamerV3 optimizes a policy in the latent space using imagination-based trajectories and actor-critic learning. Unlike ALM, which uses a single objective for optimizing both the policy and latent model, DreamerV3 separates these processes and learns a latent world model using multiple objectives. Please refer to Appendix A for further details.
@@ 3 Performance Gap of Dyna-based MBRL Across Benchmarks
In this section, we put forth the following hypothesis: incorporating Dyna-style modifications into otherwise successful off-policy algorithms can prevent them from improving beyond their performance at initialization — that of a randomly initialized policy. In support of this hypothesis we present empirical evaluations of MBPO’s performance across six OpenAI Gym tasks and six challenging robotics tasks (Nikishin et al., 2022) from DMC using our JAX-based (Bradbury et al., 2018) implementation of MBPO. For the OpenAI Gym benchmark, we demonstrate that our implementation matches the performance produced by the original PyTorch (Paszke et al., 2019) MBPO implementation (Janner et al., 2019) and ALM (Ghugare et al., 2022). Furthermore, we show that MBPO dramatically outperforms its “no Dyna” base off-policy algorithm in OpenAI Gym, Soft Actor-Critic (SAC) (Haarnoja et al., 2019a), even when both share identical hyperparameters and architectures.
In contrast, in DMC (Tassa et al., 2020) we observe that MBPO struggles to achieve any policy improvements in six out of fifteen [^1] Training curves for MBPO in all fifteen challenging DMC environments are provided in Appendix G in Figure 11. of the challenging environments that we tested on when training from scratch. Because MBPO differs from SAC only due to Dyna-style “enhancements,” these results demand a deeper investigation into whether the default SAC hyperparameters, the model ensemble, or other factors induce an otherwise healthy SAC implementation to fail.
This performance gap extends beyond MBPO into other model-based RL algorithms in the Dyna family (Sutton, 1991). To this end, we deploy the recently developed ALM (Ghugare et al., 2022) method across the same six [^2] Training curves for ALM in all fifteen challenging DMC environments are provided in Appendix G in Figure 17. challenging DMC tasks and observe the same phenomenon.
@@ 3.1 Results in OpenAI Gym
Here, 1-step and $n$-step refer to the length of synthetic rollouts simulated using the learned world model. It is common to prefer shorter rollouts, however, in order to mitigate compounding model errors (Janner et al., 2019; Sikchi et al., 2022); indeed, we demonstrate that these errors are quite substantial in Section 4.1. Combining the results in Figure 2 and Figure 3 we can conclude that our JAX-based implementation of 1-step MBPO matches the original implementation’s (Janner, 2019) $n$-step MBPO’s performance with much greater wall-clock speed, and as such we will henceforth refer to 1-step MBPO as simply “MBPO.”
Figure 3: 1-step MBPO consistently achieves higher final episodic returns and demonstrates faster convergence compared to SAC. Likewise, 1-step MBPO either matches or surpasses ALM in terms of sample efficiency across all but one task. Solid curves correspond to the mean and shaded regions represent the standard deviation across six trials. The dotted line represents the episodic return of the original $n$-step MBPO at the end of training.
@@ 3.2 Results in the DeepMind Control Suite
In contrast, MBPO’s performance when training from scratch drops significantly in the DMC environments despite using the exact same hyperparameters and model structure as for the successes in OpenAI Gym. Full results across all fifteen challenging DMC environments are provided in Figure 11 in Appendix G. These results show the complete range of MBPO’s performance, including scenarios where MBPO solves tasks more slowly than SAC, matches SAC’s sample efficiency, or fails to make any policy improvements.
@@ 3.3 Beyond MBPO: Another Dyna-based Algorithm
In this section, we show that the conclusions in Section 3.2 extend beyond MBPO to another member of the Dyna family, Aligned Latent Models (ALM) (Ghugare et al., 2022). We choose ALM for two reasons. First, unlike MBPO, ALM is built on top of DDPG, which allows us to test Dyna in a deterministic policy framework, in contrast to SAC’s stochastic approach. Second, DDPG’s reliance on Gaussian action noising offers a distinct exploration strategy compared to SAC’s entropy-based exploration.
To evaluate whether Dyna-based “enhancements” negatively impact ALM, we repeated the same experiment as for MBPO across six [^3] Training curves for ALM in all fifteen challenging DMC environments are provided in Appendix G in Figure 17. challenging DMC tasks. We train ALM with and without Dyna-style enhancements using identical hyperparameters, and compare their performance. This “no-Dyna” variant of ALM is inspired by (Ghugare et al., 2022) where the authors replaced actor training via synthetic model rollouts with a TD3 (Fujimoto et al., 2018) actor loss that relies only on real transitions.
@@
Figure 5: Performance of ALM with and without Dyna-style enhancements on six DMC benchmarks. Like MBPO, ALM shows no policy improvement with Dyna-style enhancements, but performs strongly without them (6 trials, $\pm$ std).
@@ 4.1 Is High Model Error the Problem?
In Section 3, we saw that Dyna-style “enhancements” caused both MBPO and ALM to dramatically underperform their “no Dyna” counterparts in DMC, whereas the same enhancements led to performance gains in OpenAI Gym. Since the predictive model ensemble is central to MBPO’s sample efficiency gains over model-free methods, we focus on this first as a source of MBPO’s issues. Studies have shown that high model error causes significant performance degradation (Gu et al., 2016; Rajeswaran et al., 2017), and models are most effective over short to moderate planning horizons (Sikchi et al., 2022; Holland et al., 2019) if compounding model errors can be mitigated (Talvitie, 2014; Buckman et al., 2018). We find that severe model error in many DMC environments makes even 1-step rollouts unrealistic, rendering multi-step methods unproductive.
To quantify this effect, we introduce the percent model error, defined as:
$\frac{\|\hat{y}-y\|_{2}}{\|y\|_{2}}\times 100,$
where $\hat{y}$ denotes the model’s predicted next observation and reward, and $y$ is the corresponding ground-truth target. This normalized metric allows for fair comparison across tasks with differing scales.
@@
The relationship between percent model error and performance degradation in MBPO is challenging to assess, particularly within DMC environments. We investigated this via comparisons of percent model errors across the training distribution (i.e., the replay buffer) for six tasks in both OpenAI Gym and DMC (cf. Figures 12 and 13 in Appendix G) instead of relying on computationally expensive on-policy rollouts. We observed that in all six OpenAI Gym tasks the model error converges significantly below $25\%$. In contrast, for the six DMC tasks where MBPO fails, the tasks fall into two categories. The hopper tasks exhibit model errors that converge above $100\%$, while the remaining tasks have model errors converging above $25\%$ as training progresses. Notably, all of these environments fail to improve policy performance beyond that of a random policy (recall Figure 4).
@@
While modeling errors in MBPO for DMC environments are not universally larger than those in OpenAI Gym, their impact is difficult to contextualize without further analysis. To investigate this impact, and inspired by previous work (Kalweit & Boedecker, 2017), we vary the synthetic-to-real data ratio, $S$, which represents the proportion of synthetic rollouts in each training batch. By varying $S$, we can evaluate a spectrum between MBPO ($S\to 0.95$) and SAC ($S\to 0$), thereby systematically studying the effect of synthetic data on policy performance and assessing the significance of modeling errors in degrading MBPO’s performance (Gu et al., 2016).
Focusing on hopper-stand and humanoid-stand, which represent the two categories of model error magnitude and highlight where MBPO consistently fails with a learned model, we find a clear trend: increasing $S$ reduces episodic returns (cf. Figure 14 in Appendix G). This effect is particularly evident for humanoid-stand, where even minimal synthetic data severely degrades performance, highlighting the detrimental effect of model errors on MBPO.
@@
To rule out tuning issues, we performed comprehensive hyperparameter sweeps for hopper-stand and humanoid-stand (cf. Appendix F), since introducing any synthetic data can induce failures in MBPO with the default hyperparameters. We swept over the key hyperparameters related to the training and utilization of the predictive model, including the size of the model’s hidden layers, the interval at which the model is retrained, the learning rate, and the number of training steps. None provided any improvement to policy performance across training despite reductions in model-error, which suggests that deeper architectural and algorithmic changes are required. Reduction of model bias remains an open issue in the broader MBRL literature (Wang et al., 2019; Deisenroth & Rasmussen, 2011; Luo et al., 2024) and we aim to address this in future work as part of a broader solution to Dyna-style algorithm issues identified here.